OpenStudy (anonymous):

find the domain and range..x=(y-1)^2-2??

OpenStudy (anonymous):

Domain: x>= -2 Range: all real y

OpenStudy (anonymous):

how did you find that?..i am a little lost, need a little more explanation..

OpenStudy (anonymous):

First you solve for y: \[x=\left( y-1 \right)^{2}-2\] \[x+2=\left( y-1 \right)^{2}\] \[\sqrt{x+2}=y-1\] \[y=1+\sqrt{x+2}\] We can only have positive values under the square root sign, or our answer is no longer in the real numbers. \[x+2\ge0\] So our domain is: \[x\ge-2\] Now, for the range, our output from the square root sign is never going to be negative. So the minimum value we can have for y is 1 (when x=0). So our range is: \[y \ge1\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhh

OpenStudy (anonymous):

u sure mooby ?

OpenStudy (anonymous):

when you take sqrt you get +-

OpenStudy (anonymous):

\[y= 1+\sqrt{x+2} \] is not the same as \[x= (y-1)^2 -2 \]

OpenStudy (anonymous):

range is all y